OpenStudy (anonymous):

Purnima filled her garden with 8 cubic feet of organic dirt. The garden is a rectangular prism with a volume V equals l times w times h where l is length, w is width, and h is depth. If the garden has a depth of 0.8 feet and is 2 feet wide, how long is the garden? 5 feet 2 feet 12.8 feet 3.2 feet

OpenStudy (mertsj):

\[V=l w h\]

OpenStudy (mertsj):

Replace V with 8 Replace h with 0.8 Replace w with 2 and solve for l

OpenStudy (praxer):

8/0.8*2=L => L = 8/1.6 =5feet